# Accreditation Against Limited Adversarial Noise

### v1

@@ III.1 Trap and Target Circuits
In this paper, I do not propose to develop new trap circuits or target circuits. In fact, I would prefer to make minimal changes to pre-existing protocols. To this end, I will talk in generalities about traps and targets, and only specify conditions on target circuits and trap circuits - met by all target and trap circuits in the accreditation protocols in Ref. [25], Ref. [26], and Ref. [27] - that they must meet. I will assume that I have two efficient classical algorithms, $P_{\textit{targ}}$ and $P_{\textit{trap}}$, that, if given a valid quantum circuit as input, return the corresponding target and a matching trap, respectively.
@@
    In target circuits and trap circuits all error occurring is twirled (via Pauli twirls or some other twirl) to stochastic error.
@@
    Each trap acts as a Bernoulli trial with the success probability being the probability error occurs in that trap according to the particular CPTP maps - representing the error - in the trap.
@@
The above assumptions have mentioned twirling CPTP error to stochastic error, which I define formally in Def. 4. This is an important step in all accreditation protocols, as it reduces the error to a known, more easily quantified form.
@@
CPTP error within a quantum circuit is said to be twirled to stochastic error if, via the addition of only single-qubit gates to the circuit, the error is effectively transformed to stochastic error, without otherwise affecting the outputs of the circuit. Likewise, gates are said to be twirled if any error occurring in them is twirled. An important example is Pauli twirling [49, 50, 51], where the resulting stochastic error consists exclusively of Pauli operators.
My new adversarial error model may, however, require a slight modification of the trap circuits and target circuits: the error model they are designed for may not consider “hiding” the outputs as there was no adversary to “hide” them from. With my new adversarial error model, this becomes necessary as otherwise Bob may be able to identify traps and targets. The required hiding / encrypting of outputs is achieved via Algorithm 2.
@@
I note that the required single-qubit unitaries, $\mathcal{U}_{M}$, will exist for any single-qubit measurement and that the outcome of the circuit - if the single-qubit gates of the circuit are not known, as is the case for Bob as they are redacted for him - is irretrievable from the measurement outcomes without the key. This hiding of the circuit outputs is information-theoretically secure (i.e. has perfect security) [52], assuming the single-qubit gates are redacted, and comparable to the one-time pads in universal blind quantum computing [53]. Henceforth, I will denote $P_{\textit{targ}}$ and $P_{\textit{targ}}$, with the changes presented above to make their outputs hidden, by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$ respectively.
A similar method may be used to hide the inputs into circuits.
@@ III.2 Presentation of the Upgraded Accreditation Protocol
With trap and target circuits established, I can consider the full accreditation protocol to be developed. In line with Sec. III.1, for Sec. III.2 I assume I already have some method of generating trap and target circuits conforming to the requirements in Sec. III.1. However, before the full protocol can be presented, in Protocol 3, I must establish the statistical foundations of the new protocol - in Sec. III.2.1 - and the core mechanics of the protocol - in Sec. III.2.2.
@@ III.2.1 Statistical Basis of the New Accreditation Protocol
Before presenting the accreditation protocol, it is first useful to present a purely statistical lemma (Lemma 2) that will later enable the accreditation protocol. This presentation first begins with Def. 5 and Lemma 1.
@@
I may then present and prove Lemma 2, which makes use of Lemma 1. Whereas Lemma 1 was concerned with characterizing the interval $\mathcal{R}$, Lemma 2 aims to place a bound on any value occurring within $\mathcal{R}$.
@@ III.2.2 Core Mechanics of the New Accreditation Protocol
Before its formal presentation, in Sec. III.2.3, the core mechanics of the accreditation protocol are presented. This takes the form of Lemma 3 and Lemma 4.
@@
If given a classical polynomial-time algorithm, $P^{\prime}_{\textit{trap}}$, generating traps corresponding to a circuit, $\mathcal{C}$, and the corresponding $m$, and $k$ (as defined in Sec. III.1); assuming the probability of error in each execution of a trap is independent of the outcomes of all preceding traps, that any trap detects any specific error with probability at least $k\in[0,1]$, and all traps are executed via Protocol 1, $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps are sufficient to estimate, $\mathrm{Avg}(\mathcal{R})$ (where $\mathcal{R}$ is the set of probabilities of error in each trap), to within error $\theta$ with confidence $\alpha$.
Proof.
Due to the independence, between traps, of whether error occurs, as assumed, $\mathrm{Avg}(\mathcal{R})$ can be estimated, as required, using Hoeffding’s inequality [54] and $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps, as the fraction of traps returning an incorrect measurement outcome serves as an approximation to the probability of the traps not measuring the correct results, to required error. This then provides the required bound - to required error - by dividing it by $k$. ∎
With Lemma 3 established, I can present and prove a central theorem of the the accreditation protocol’s core mechanics: Lemma 4.
Lemma 4.
If given a polynomial algorithm, $P_{\textit{trap}}^{\prime}$, generating traps for a fixed circuit, $\mathcal{C}$, and the corresponding $m$, and $k$ (as defined in Sec. III.1); assuming the probability of error in each execution is independent of the outcomes of all preceding executions, that any trap detects any specific error with probability at least $k\in[0,1]$, and all traps are executed via Protocol 1, the probability of error in any execution of any circuit executed within the same singular use of Protocol 1 - such as a target circuit  [^6] As the probability of error in the target would, due to the assumptions made, be within $[(1-\beta)P_{0},(1+\beta)P_{0}]$. - as the traps can be upper bounded by:
@@
As in Lemma 3, define the set $\mathcal{R}$ to be the probability of error occurring in each specific trap in a set of traps executed, along with any number of other circuits (that are assumed to meet the conditions to be executed alongside the traps in Protocol 1), in a singular use of Protocol 1. This implies that the probability of error in any specific trap mentioned above - and in any circuit executed within the same single use of Protocol 1 - is within the interval:
@@
The probability a specific trap returns an incorrect measurement outcome, given error occurs, is, by assumption of having the required traps, lower bounded by $k\in[0,1]$; hence, an upper bound on $\mathrm{Avg}(\mathcal{R})$ can be approximated, as in Lemma 3, to within error $\theta$, with confidence $\alpha$. Let $\bar{\nu}$ denote the approximation achieved from sampling using the traps (i.e. $\bar{\nu}$ is the fraction of traps measuring in incorrect output). Then, with confidence, $\alpha$,
@@
As per Lemma 3, which is applicable due to the assumption that the probability of error is independent of previous measurement outcomes - this requires only $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps. ∎
III.2.3 Formal Protocol For Accreditation
The core components of my accreditation protocol have now been constructed and presented - in Sec. III.2.1 and Sec. III.2.2 - so I can now present my accreditation protocol formally, in Algorithm 3. This algorithm is presented as would be used by Alice, presupposing the error model in Sec. II.2.
@@ III.2.3 Formal Protocol For Accreditation
$\bullet$ A circuit, $\mathcal{C}$, meeting any required conditions.
@@
Algorithm 3 Formal Accreditation Protocol
@@
Assuming all sets of circuits executed are executed via Protocol 1, and that the classical algorithms, $P^{\prime}_{\textit{targ}}$ and $P^{\prime}_{\textit{trap}}$, generate trap and target circuits conforming to the conditions in Sec. III.1, then Algorithm 3 provides a sample from a probability distribution and a bound on the ideal-actual variation distance between that distribution and the distribution generated by the measurements in $\mathcal{C}$ when no error occurs (to within accuracy $\theta$, with confidence $\alpha$).
@@
where $\bar{v}$ is the fraction of traps returning an incorrect measurement result and $k\in[0,1]$ is as defined in Sec. III.1. As mentioned, for Eqn. 16 to hold, it is required that the probability of error in each trap is independent of the outcomes of every prior trap. This is not initially guaranteed as Bob is free to choose probabilities based on previous measurement outcomes. However, due to the changes made to each trap and target in Algorithm 2, to hide the outputs of each trap from Bob with perfect security, Bob does not know the result of any trap or target. Therefore, he cannot make any of his choices based on the results of any prior trap or target  [^7] I additionally note that Bob cannot remove this independence by basing his choice of CPTP maps on error in previous traps as although Bob knows which CPTP maps are applied to all previous circuits executed, he does not know which stochastic errors have occurred due to them, only the probabilities of each occurring (due to having chosen the CPTP maps) because he doesn’t control any aspect of the CPTP maps’ applications, due to the construction of Protocol 1. Therefore, Bob cannot base the choice of CPTP maps in a trap on the outcome of any stochastic property of previous traps as they are unknown to him.. Hence, each trap can be considered to be independent i.e. the probability of error in each trap or target is the result of Bob’s choices but must remain completely independent of the results of any prior trap or target.
@@ IV Discussion
In this paper I have presented an easily-applicable method for upgrading any pre-existing digital accreditation protocol to considering all error to be adversarial. This has necessitated the development of a new model of adversarial error, where Bob is limited according to experimentally extracted rules. These rules - that single-qubit gates are hidden from Bob and that the probability of error in similar circuits executed in quick succession have similar probabilities of error - allow any extant accreditation protocol to function almost identically when experiencing adversarial error as they do when the error is IID. Therefore, the desirable qualities of these protocols, such as not requiring excessively long circuits, or ancilla qubits, or extensive connectivity between qubits are preserved; leading to no diminution in their suitability for near-term usage.
The principle question left open is to what extent can the limits on Bob be relaxed. This would have to entail changes to Protocol 1. I believe the redaction of single-qubit gates cannot be eliminated without substantial changes to the trap, as otherwise Bob could distinguish target and trap circuits - although it is not clear trap and target circuits that are indistinguishable to Bob can be constructed without requiring another limitation on Bob. A slight improvement can be obtained using Ref. [26, Theorem 1], which allows the error in single-qubit gates to be weakly gate dependent: this could be added to my cryptographic model by allowing Bob to define error in the - still redacted - single-qubit gates that depends on the single-qubit gates (single-qubit gates would still have to be redacted to prevent Bob identifying the target) but the differing CPTP maps - for each single-qubit gate - must differ, in the diamond norm, by, at most, some small, known, value. This is a further step towards reality and the restriction on the differences between the gate-dependent error reflects the very small errors in single-qubit gates [35, Fig.5]. Finally I note that, if I take the philosophical position that error that does not affect the state of the qubits effectively does not occur and can be considered to have never been applied, I can relax the assumption that Bob does not directly control stochastic processes in the CPTP maps modelling error. This is as the Pauli twirling, accompanied by the random Hadamard gates, make it impossible for Bob to know whether gates he has applied have any affect - which, as mentioned above, can be interpreted as him not being able to decide if he has effectively applied the error. This breaks any ability of Bob to apply error in a way that is dependent on the error in previous circuit executions. Due to this, the error can be considered to be independent between circuit executions.

### v2

@@ III Accreditation Protocol
Sec. III is dedicated to resolving the problem Alice faces in the problem setting established in Sec. II.3. This is equivalent to upgrading an accreditation protocol that assumed error is CPTP and IID (and that single-qubit gates suffer only gate-independent error) to one that works with the error assumptions in Sec. II.3.
@@ III.1 Trap and Target Circuits
In this paper, I do not propose to develop new trap circuits or target circuits. In fact, I would prefer to make minimal changes to the trap and target circuits in Ref. [23]. I will also not regurgitate the exact designs of the trap and target circuits in Ref. [23] and instead - throughout this paper - will assume that I have two efficient classical algorithms, $P_{\textit{targ}}$ and $P_{\textit{trap}}$, that, if given any quantum circuit as input, return the corresponding target and a random [^9] I.e. $P_{\textit{trap}}$ will choose the random gates for the Pauli twirl and error detection and so will return a slightly different circuit each time. matching trap, respectively, of the protocol in Ref. [23].
I briefly note the important features of these trap and target circuits:
@@
    In target circuits and trap circuits all error occurring is twirled (via Pauli twirls or some other twirl) to stochastic Pauli error.
@@
    Each execution of a trap acts as a Bernoulli trial, with the success probability - for that trial - being the probability of error occurring in that trap execution.
@@
The above assumptions have mentioned twirling CPTP error to stochastic Pauli error, which I define formally in Def. 8. This is an important step in all accreditation protocols, as it reduces the error to a known, more easily quantified form.
@@
CPTP error within a quantum circuit is said to be twirled to stochastic Pauli error [45, 46, 47] if, via the addition of only single-qubit Pauli gates to the circuit, the error is effectively transformed to stochastic Pauli error, without otherwise affecting the outputs of the circuit. Likewise, gates are said to be twirled if any error occurring in them is twirled.
As herein noise is now considered to be adversarial, the accreditation protocol of this paper does require a slight modification of the trap circuits and target circuits: Ref. [23] did not consider “hiding” the measurement outcomes of a circuit as there was no adversary to “hide” them from (according to its error model). With adversarial noise (i.e. the problem setting of Sec. II.3), this becomes necessary as otherwise Bob may be able to identify traps and targets based on these outcomes, or base future error on the measurement outcomes of previous circuits. The required hiding / encrypting of measurement outcomes is achieved via Algorithm 2.
@@
I note that the required (by Algorithm 2) single-qubit unitaries, $\mathcal{U}_{M}$, will exist for any single-qubit measurement and that the outcome of the circuit - if the single-qubit gates of the circuit are not known, as is the case for Bob as they are redacted for him - is irretrievable from the measurement outcomes without the key. This hiding of the circuit outputs is information-theoretically secure (i.e. has perfect security) [48], assuming the single-qubit gates are redacted, and comparable to the one-time pads in universal blind quantum computing [49]. Henceforth, I will denote $P_{\textit{targ}}$ and $P_{\textit{targ}}$, with the changes presented above to make their outputs hidden, by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$ respectively.
@@ III.2 Presentation of the Upgraded Accreditation Protocol
With trap and target circuits established, I can consider the full accreditation protocol to be developed. In line with Sec. III.1, for Sec. III.2, traps and targets will be treated as blackboxes and I will only refer to their construction as being performed by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$. Their only relevant properties will be that traps detect any error with probability $k\in[0,1]$, all error in a trap or target simulation is reduced to effective Pauli error, and Bob cannot tell the difference between trap and target simulations due to them only differing in their single-qubit gates.
However, before the full protocol can be presented, in Protocol 3, I must establish the statistical foundations of the new protocol - in Sec. III.2.1 - and the core mechanics of the protocol - in Sec. III.2.2. These statistical methods will be used to evaulate the $P_{0}$ that defines Bob’s particular choice of a SPSCLβ using multiple trap executions (still all within a single use of Protocol 1).
@@ III.2.1 Statistical Basis of the New Accreditation Protocol
Before presenting the accreditation protocol, it is first useful to present a purely statistical lemma (Lemma 4) that will later enable the accreditation protocol. This presentation first begins with Def. 9 and Lemma 3.
@@
I may then present and prove Lemma 4, which makes use of Lemma 3. Whereas Lemma 3 was concerned with characterizing the interval $\mathcal{R}$, Lemma 4 aims to - using Lemma 3 - place a bound on any value occurring within $\mathcal{R}$ by estimating a bound on $P_{0}$.
@@ III.2.2 Core Mechanics of the New Accreditation Protocol
Before its formal presentation, in Sec. III.2.3, the core mechanics of the accreditation protocol are presented. This takes the form Lemma 5.
@@
If given a classical polynomial algorithm, $P_{\textit{trap}}^{\prime}$, generating traps corresponding to a fixed circuit, $\mathcal{C}$, and the corresponding $m$, and $k$ (as defined in Sec. III.1); assuming the probability of error in each execution is independent of the outcomes of all preceding executions, that any trap detects any specific error with probability at least $k\in[0,1]$, and all traps are executed via Protocol 1, the probability of error in any execution of any circuit executed within the same singular use of Protocol 1 - such as a target circuit  [^10] As the probability of error in the target would, due to the assumptions made, be within $[(1-\beta)P_{0},(1+\beta)P_{0}]$. - as the traps can be upper bounded by:
@@
Let $\mathcal{R}$ be an ordered set where the $j$th element is the probability of error occurring in the $j$th execution of a set of trap executions interspersed among any number of other circuits. In a singular use of Protocol 1. This implies that the probability of error in any specific trap mentioned above - and in any circuit executed within the same single use of Protocol 1 - is within the interval:
@@
The probability a specific trap returns an incorrect measurement outcome, given error occurs, is, by assumption of having the required traps, lower bounded by $k\in[0,1]$.
Therefore, by also using the independence between circuit executions of whether error occurs, $\mathrm{Avg}(\mathcal{R})$ can be approximated - to within error, $\theta$, with confidence $\alpha$ using Hoeffding’s inequality [50] and $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ trap executions.
This approximation of $\mathrm{Avg}(\mathcal{R})$ is simply the fraction of traps returning an incorrect measurement outcome divided by $k$ (to account for the cases where error is not detected). Let $\bar{\nu}$ denote this approximation. Then, with confidence, $\alpha$,
@@
III.2.3 Formal Presentation of the Accreditation Protocol and Proof of its Correctness
The core components of my accreditation protocol have now been constructed and presented - in Sec. III.2.1 and Sec. III.2.2 - so I can now present my accreditation protocol formally, in Algorithm 3. This algorithm is presented as would be used by Alice, presupposing the problem setting in Sec. II.4.
@@ III.2.3 Formal Presentation of the Accreditation Protocol and Proof of its Correctness
$\bullet$ A circuit, $\mathcal{C}$.
@@
Protocol 3 Formal Accreditation Protocol
@@
Assuming all sets of circuits executed are executed via Protocol 1, and that the classical algorithms, $P^{\prime}_{\textit{targ}}$ and $P^{\prime}_{\textit{trap}}$, generate trap and target circuits with the features listed in Sec. III.1 (i.e. as in Ref. [23]), then Algorithm 3 provides a sample from a probability distribution and a bound on the ideal-actual variation distance between that distribution and the distribution generated by the measurements in $\mathcal{C}$ when no error occurs (to within accuracy $\theta$, with confidence $\alpha$).
@@
where $\bar{v}$ is the fraction of traps returning an incorrect measurement result and $k\in[0,1]$ is as defined in Sec. III.1. As mentioned, for Eqn. 19 to hold, it is required that the probability of error in each trap is independent of the outcomes of every prior trap. This is not initially guaranteed as Bob is free to choose probabilities based on previous measurement outcomes. However, due to the changes made to each trap and target in Algorithm 2, to hide the outputs of each trap from Bob with perfect security, Bob does not know the result of any trap or target. Therefore, he cannot make any of his choices based on the results of any prior trap or target  [^11] I additionally note that Bob cannot remove this independence by basing his choice of CPTP maps on error in previous traps. Although Bob knows which CPTP maps are applied to all previous circuits executed, he does not know which errors have occurred due to the stochastic processes involved in the CPTP maps he chose, only the probabilities of each occurring as he doesn’t control any aspect of the CPTP maps’ applications, due to the construction of Protocol 1. Therefore, Bob cannot base the choice of CPTP maps in a trap on the outcome of any stochastic property of previous traps as they are unknown to him.. Hence, each trap can be considered to be independent i.e. the probability of error in each trap or target is the result of Bob’s choices but must remain completely independent of the results of any prior trap or target.
@@
Theorem 1 is a proof of the correctness of the protocol developed herein, but the main aim of this paper is to provide Alice with a solution to her predicament outlined in Sec. II.1 and Sec. II.3. To this end, I present another theorem - without proof - that follows directly from Theorem 1 catering to Alice more directly.
Theorem 2.
Protocol 3 allows Alice to get Bob to implement any quantum circuit of her choosing and be able to place a bound on the ideal-actual variation distance (with arbitrary confidence $\alpha$) of the results she receives.
@@ IV Discussion
In this paper I have upgraded the accreditation protocol of Ref. [23] to consider all error to be adversarial. This has necessitated the development of a new model of adversarial error, where Bob is limited according to experimentally extracted rules. These rules - that single-qubit gates are hidden from Bob and that the probability of error in similar circuits executed in quick succession have similar probabilities of error - allow my upgraded accreditation protocol to function almost identically when experiencing adversarial error as the one in Ref. [22] does when the error is IID. Therefore, the desirable qualities the protocol from Ref. [22], such as not requiring lengthening the circuit to accredit, or adding excessively many extra single-qubit gates, or any extra two-qubit gates, or ancilla qubits, or trusting any aspect of a computation, or extra connectivity between qubits are preserved; leading to no diminution in its suitability for near-term usage.
The principle question left open is to what extent can the limits on Bob be relaxed. Relaxing the limitations would have to entail changes to Protocol 1. I believe the redaction of single-qubit gates cannot be eliminated without substantial changes to the trap, as otherwise Bob could distinguish target and trap circuits - although it is not clear trap and target circuits that are indistinguishable to Bob can be constructed without requiring another limitation on Bob.
A slight improvement can be obtained using Ref. [23, Theorem 1], which allows the error in single-qubit gates to be weakly gate dependent: this could be added to my cryptographic model by allowing Bob to define error in the - still redacted - single-qubit gates that depends on the single-qubit gates  [^12] Single-qubit gates would still have to be redacted to prevent Bob identifying the target but he could - perhaps - specify some dependence without seeing the single-qubit gates e.g. “if that gate is an $\hat{X}$ gate…” but the differing CPTP maps - for each single-qubit gate - must differ, in the diamond norm, by at most some small known value. This is a further step towards reality and the restriction on the differences between the gate-dependent error reflects the very small errors in single-qubit gates [32, Fig.5].
Finally I note that, if I take the philosophical position that error that does not affect the state of the qubits, or only adds an overall phase, effectively does not occur and can be considered to have never been applied, I can relax the assumption that Bob does not directly control stochastic processes in the CPTP maps modelling error. This is as the Pauli twirling, accompanied by the random Hadamard gates, make it impossible for Bob to know whether gates he has applied have any affect - which, as mentioned above, can be interpreted as him not being able to decide if he has effectively applied the error. This breaks any ability of Bob to apply error in a way that is dependent on the error in previous circuit executions. Due to this, the error can be considered to be independent between circuit executions.